OpenStudy (anonymous):

6. The area of a rectangular field is given by the trinomial t2 – 4t – 45. The length of the rectangle is t + 5. What is the expression for the width of the field?

Directrix (directrix):

t-9

OpenStudy (hoblos):

t2 – 4t – 45 =(t+5)(t-9) (t2 – 4t – 45)/(t+5) = t-9

Directrix (directrix):

A = lw t^2 -4t - 45 = (t+5)(t-9) --> width is (t-9) Questions?